Weekly revenues for a particular curry puff seller in Ang Mo Kio are normally distributed with mean $11,480 and standard deviation $1,165. For how many weeks per year would you expect revenues to be less than $8,150?


1
Expert's answer
2021-11-04T12:25:13-0400

"P(X<8150)=P(\frac{x-\mu}{\sigma}<\frac{8150-11480}{1165})"


"=P(Z<-2.86)"


From Z normal distribution table;


"P(X<8150)=0.0021"


This means only 0.21% weeks per year would expect revenues to be less than $8,150.